OpenStudy (anonymous):

In the equation, 81=3^x (the exponent is x) How would you solve this?

OpenStudy (anonymous):

81 = 3^4

OpenStudy (anonymous):

3^4 = 3^x x= 4

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

Do u Know Prime Factoriztaion ?

OpenStudy (jiteshmeghwal9):

a^b=a^c <=> b=c remember this property :)